### v1

@@ 5 Full Implementation
For all $\delta>0$, there exists $\alpha>0$ such that for all $\widehat{\boldsymbol{y}}$, $\Phi_{\pi}(\widehat{\boldsymbol{y}})\leq\min\Phi_{\pi}+\alpha$ implies that for all $a$ and all $\boldsymbol{\tau}$, $\left\lvert y_{a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert\leq\delta$.
Proof of 5.6.
By contradiction, suppose that there exists $\delta>0$ such that for all $\alpha=1/n$ with $n\in\mathbb{N}$, there exists $\widehat{\boldsymbol{y}}_{n}$ with $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{n})\leq\min\Phi_{\pi}+1/n$ and $\max_{a,\boldsymbol{\tau}}\left\lvert y_{n,a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert>\delta$. By compactness, take a subsequence and assume that $\widehat{\boldsymbol{y}}_{n}\to\widehat{\boldsymbol{y}}_{\infty}$ as $n\to\infty$. Letting $n\to\infty$, we get that $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{\infty})\leq\min\Phi_{\pi}$ and $\max_{a,\boldsymbol{\tau}}\left\lvert\widehat{y}_{\infty,a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert>\delta$. This contradicts the uniqueness of $y^{*}_{a}(\boldsymbol{\tau})$. ∎
Take a Bayesian Wardrop $\alpha$-equilibrium with flow vector $(\tilde{\boldsymbol{y}}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ and outcome $\tilde{\mu}$. We have for each $\theta\in\Theta$,
$\begin{split}\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\tilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert&=\left\lvert\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\psi(\boldsymbol{y}^{*}(\boldsymbol{\tau}),\theta)-\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\psi(\tilde{\boldsymbol{y}}(\boldsymbol{\tau}),\theta)\right\rvert\\ &\leq\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\left\lvert\psi(\boldsymbol{y}^{*}(\boldsymbol{\tau}),\theta)-\psi(\tilde{\boldsymbol{y}}(\boldsymbol{\tau}),\theta)\right\rvert\\ &\leq\omega_{\psi}(\delta).\end{split}$
@@
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\tilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\omega_{\psi}(\delta).$ (5.16)
@@
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\tilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha+2\omega_{\psi}(\delta),$ (5.17)
which is less or equal to $\varepsilon$ for $\alpha,\delta$ small enough. ∎
Proof of Theorem 5.3.
The result follows from the fact that the unique equilibrium outcome $\mu^{\ast}$ is an Bayesian Wardrop $\alpha$-equilibrium outcome for any $\alpha>0$. As in the proof of Theorem 4.1, if $\mu$ has finite support with rational flows, then we can choose $\varepsilon=0$, and the obedient outcome $\breve{\mu}_{\mathcal{I}}$ is the unique Bayesian Wardrop equilibrium outcome $\mu^{\ast}$. ∎
The next theorem shows that in games with weakly convex potential, we can fully implement the ex ante expected social cost of any Bayes correlated Wardrop equilibrium, where the social cost in state $\theta$ given flows $\boldsymbol{y}$ is:
$\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)=\sum_{a}y_{a}c_{a}(\boldsymbol{y},\theta).$ (5.18)
@@
Let $\Gamma$ be a basic game that has a convex potential in each state and let $\mu$ be a BCWE. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that for all Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$,
$\left\lvert\sum_{\theta}p(\theta)\int\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\sum_{\theta}p(\theta)\int\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon.$ (5.19)
@@
Theorem 5.7 applies to all congestion games with increasing costs. For instance, in Example 1, it implies that there exists an information structure (the one described on page 1) that implements a social cost-minimizing BCWE in every Bayesian Wardrop equilibrium of the basic game extended with this information structure. However, Theorem 5.7 does not apply to Example 2, which is a potential game but lacks convexity. In this case, Theorem 4.1 implies that there exists an information structure (the one described on page 2) that implements the social cost-minimizing BCWE in a Bayesian Wardrop equilibrium of the basic game extended with this information structure, but other, less efficient equilibria also exist. [^7] Theorem 5.7 also does not apply to games with positive payoff externalities (i.e., strictly decreasing costs), such as those in Li et al., (2023). However, Li et al., (2023) provide an example showing that full implementation may still be possible with complex information structures involving an infinite set of signals.
Proof of Theorem 5.7.
@@
Assume that $\mu$ is a BCWE with finite support, choose the function $\psi$ as the social cost function $\operatorname{\mathsf{SC}}$, and fix $\alpha>0$. Consider the information structure $\mathcal{I}$ constructed in the proof of Theorem 4.1 with parameter $\eta>0$, such that the obedient flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ with outcome $\breve{\mu}_{\mathcal{I}}$ is a Bayesian Wardrop $\alpha$-equilibrium and
@@
$\left\lvert\int\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha,\quad\forall\theta\in\Theta.$ (5.20)
@@
We know that the set of Wardrop equilibria of the associated multi-population game with complete information is the set of minimizers of the potential (Sandholm,, 2001, Proposition 3.1). Since there is a convex potential in each state, the potential defined in Eq. 5.4 is a convex function and the set of Wardrop equilibria is convex. Also, from Koessler et al., (2024, Proposition 4), all Wardrop equilibria have the same cost profile. That is, for any Wardrop equilibrium flow vectors $\boldsymbol{y}^{*}=(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ and $\boldsymbol{y}^{**}=(\boldsymbol{y}^{**}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$, for every $k,\tau^{k}$,
$C_{a^{*}}^{k,\tau^{k}}(\boldsymbol{y}^{*})=C_{a^{**}}^{k,\tau^{k}}(\boldsymbol{y}^{**}),$ (5.21)
@@
where $a^{*}$ (resp. $a^{**}$) is any action in the support of $\boldsymbol{y}^{*}$ (resp. $\boldsymbol{y}^{**}$) for population $(k,\tau^{k})$. It follows that all Wardrop equilibria have the same total cost, where the total cost of an interim flow profile $\widehat{\boldsymbol{y}}$ is,
@@
$\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}})\coloneqq\sum_{k,\tau^{k}}\sum_{a}y_{a}^{k}(\tau^{k})C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}).$ (5.22)
@@
For any interim flow profile $\widehat{\boldsymbol{y}}$, the total cost is the expectation of the social cost function.
@@
From (5.3) we have that the total cost in the population game is
$\begin{split}\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}})&=\sum_{k,\tau^{k}}\sum_{a}y_{a}^{k}(\tau^{k})\sum_{\theta,\boldsymbol{\tau}^{-k}}p(\theta)\pi\left(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta\right)c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\sum_{a}\sum_{k}y_{a}^{k}(\tau^{k})c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\sum_{a}y_{a}(\boldsymbol{\tau})c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta).\end{split}$ (5.23)
Corollary 5.9.
The expected social cost is the same in all Bayesian Wardrop equilibrium outcomes of $\Gamma$ with information structure $\mathcal{I}$.
From 5.5, any Bayesian Wardrop $\alpha$-equilibrium of the game with information structure $\mathcal{I}$, is a $\alpha$-minimizer of the associated potential. To complete the proof, we use the following argument:
Claim 5.10.
For all $\delta>0$, there exists $\alpha>0$ such that for all $\widehat{\boldsymbol{y}}$, $\Phi_{\pi}(\widehat{\boldsymbol{y}})\leq\min\Phi_{\pi}+\alpha$ implies
$\left\lvert\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}})-\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}}^{*})\right\rvert\leq\delta,$ (5.24)
Proof of 5.10.
By contradiction, suppose that there exists $\delta>0$ such that for all $\alpha=1/n$ with $n\in\mathbb{N}$, there exists $\widehat{\boldsymbol{y}}_{n}$ with $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{n})\leq\min\Phi_{\pi}+1/n$ and $\left\lvert\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}}_{n})-\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}}^{*})\right\rvert>\delta$. By compactness, take a subsequence and assume that $\widehat{\boldsymbol{y}}_{n}\to\widehat{\boldsymbol{y}}_{\infty}$ as $n\to\infty$. Letting $n\to\infty$, we get that $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{\infty})\leq\min\Phi_{\pi}$ and $\left\lvert\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}}_{\infty})-\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}}^{*})\right\rvert>\delta$. This is a contradiction since $\widehat{\boldsymbol{y}}_{\infty}$ would then be a minimizer of the potential, thus a Bayesian Wardrop equilibrium, with a different total cost. ∎
This concludes the proof of Theorem 5.7. ∎
6 List of Symbols
$\mathcal{A}$ set of action profiles, defined in Definition 2.1

### v2

@@ 5.2 Full Implementation
For all $\delta>0$, there exists $\alpha>0$ such that for all $\widehat{\boldsymbol{y}}$, $\Phi_{\pi}(\widehat{\boldsymbol{y}})\geq\max\Phi_{\pi}-\alpha$ implies that for all $a$ and all $\boldsymbol{\tau}$, $\left\lvert y_{a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert\leq\delta$.
Proof of ˜5.8.
By contradiction, suppose that there exists $\delta>0$ such that for all $\alpha=1/n$ with $n\in\mathbb{N}$, there exists $\widehat{\boldsymbol{y}}_{n}$ with $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{n})\geq\max\Phi_{\pi}-1/n$ and $\max_{a,\boldsymbol{\tau}}\left\lvert y_{n,a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert>\delta$. By compactness, take a subsequence and assume that $\widehat{\boldsymbol{y}}_{n}\to\widehat{\boldsymbol{y}}_{\infty}$ as $n\to\infty$. Letting $n\to\infty$, we get that $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{\infty})\geq\max\Phi_{\pi}$ and $\max_{a,\boldsymbol{\tau}}\left\lvert\widehat{y}_{\infty,a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert\geq\delta$. This contradicts the uniqueness of $y^{*}_{a}(\boldsymbol{\tau})$. ∎
Take a Bayesian Wardrop $\alpha$-equilibrium with flow vector $(\widetilde{\boldsymbol{y}}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ and outcome $\widetilde{\mu}$. We have for each $\theta\in\Theta$,
$\begin{split}\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\widetilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert&=\left\lvert\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\psi(\boldsymbol{y}^{*}(\boldsymbol{\tau}),\theta)-\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\psi(\widetilde{\boldsymbol{y}}(\boldsymbol{\tau}),\theta)\right\rvert\\ &\leq\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\left\lvert\psi(\boldsymbol{y}^{*}(\boldsymbol{\tau}),\theta)-\psi(\widetilde{\boldsymbol{y}}(\boldsymbol{\tau}),\theta)\right\rvert\\ &\leq\omega_{\psi}(\delta).\end{split}$
@@
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\widetilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\omega_{\psi}(\delta).$ (5.26)
@@
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\widetilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha+2\omega_{\psi}(\delta),$ (5.27)
which is smaller than $\varepsilon$ for $\alpha,\delta$ small enough. ∎
Proof of Theorem˜5.5.
The result follows from the fact that the unique equilibrium outcome $\mu^{\ast}$ is an Bayesian Wardrop $\alpha$-equilibrium outcome for any $\alpha>0$. As in the proof of Theorem˜4.1, if $\mu$ has finite support with rational flows, then we can choose $\varepsilon=0$, and the obedient outcome $\breve{\mu}_{\mathcal{I}}$ is the unique Bayesian Wardrop equilibrium outcome $\mu^{\ast}$. ∎
The next theorem shows that in games with a weakly concave potential we can fully implement the ex-ante expected total payoff of any Bayes correlated Wardrop equilibrium, where the total payoff in state $\theta$ given flows $\boldsymbol{y}$ is:
$\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)\coloneqq\sum_{a}y_{a}u_{a}(\boldsymbol{y},\theta).$ (5.28)
@@
Let $\Gamma$ be a basic game with a concave potential in each state and let $\mu$ be a BCWE. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that for each Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$,
$\left\lvert\sum_{\theta}p(\theta)\int\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\sum_{\theta}p(\theta)\int\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon.$ (5.29)
@@
Theorem˜5.9 applies to Cournot games. In Example˜3.2, it implies that there exists an information structure (the one described on page 3.2) that implements a BCWE which maximizes total profit in every Bayesian Wardrop equilibrium of the basic game extended with this information structure. However, Theorem˜5.9 does not apply to Example˜3.1, which is a potential game but lacks concavity. In this case, Theorem˜4.1 implies that there exists an information structure (the one described on page 3.4) that implements the BCWE which maximizes the total payoff, in a Bayesian Wardrop equilibrium of the basic game extended with this information structure, but other, less efficient equilibria also exist. [^7] Theorem 5.9 does not apply to games with positive payoff externalities such as those in Li et al., (2023). However, Li et al., (2023) provide an example showing that full implementation may still be possible with complex information structures involving an infinite set of signals.
Proof of Theorem˜5.9.
@@
Assume that $\mu$ is a BCWE with finite support, choose the function $\psi$ as the social payoff function $\operatorname{\mathsf{U}}$, and fix $\alpha>0$. Consider the information structure $\mathcal{I}$ constructed in the proof of Theorem˜4.1 with parameter $\eta>0$, such that the obedient flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ with outcome $\breve{\mu}_{\mathcal{I}}$ is a Bayesian Wardrop $\alpha$-equilibrium and
@@
$\left\lvert\int\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha,\quad\forall\theta\in\Theta.$ (5.30)
@@
We know that the set of Wardrop equilibria of the associated multi-population game with complete information is the set of maximizers of the potential (Sandholm,, 2001, Proposition 3.1). Because there is a concave potential in each state, the potential defined in Eq.˜5.14 is a concave function and the set of Wardrop equilibria is convex. Also, from Koessler et al., (2025, Proposition 4), all Wardrop equilibria have the same payoff profile. That is, for any Wardrop equilibrium flow vectors $\boldsymbol{y}^{*}=(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ and $\boldsymbol{y}^{**}=(\boldsymbol{y}^{**}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$, for every $k,\tau^{k}$,
$U_{a^{*}}^{k,\tau^{k}}(\boldsymbol{y}^{*})=U_{a^{**}}^{k,\tau^{k}}(\boldsymbol{y}^{**}),$ (5.31)
@@
where $a^{*}$ (resp. $a^{**}$) is any action in the support of $\boldsymbol{y}^{*}$ (resp. $\boldsymbol{y}^{**}$) for population $(k,\tau^{k})$. It follows that all Wardrop equilibria have the same total payoff, where the total payoff of an interim flow profile $\widehat{\boldsymbol{y}}$ is,
[…]
From (5.13), the total payoff in the population game is equal to the expected total payoff in the basic game, i.e.,
@@
$\begin{split}\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}})&=\sum_{k,\tau^{k}}\sum_{a}y_{a}^{k}(\tau^{k})\sum_{\theta,\boldsymbol{\tau}^{-k}}p(\theta)\pi\left(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta\right)u_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\sum_{a}\sum_{k}y_{a}^{k}(\tau^{k})u_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\sum_{a}y_{a}(\boldsymbol{\tau})u_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\operatorname{\mathsf{U}}(\boldsymbol{y},\theta).\end{split}$ (5.33)
It follows that the expected total payoff is the same in all Bayesian Wardrop equilibrium outcomes of $\Gamma$ with information structure $\mathcal{I}$. From ˜5.7, any Bayesian Wardrop $\alpha$-equilibrium of the game with information structure $\mathcal{I}$, is a $\alpha$-maximizer of the associated potential. To complete the proof, we use the following argument:
@@
For all $\delta>0$, there exists $\alpha>0$ such that for all $\widehat{\boldsymbol{y}}$, $\Phi_{\pi}(\widehat{\boldsymbol{y}})\geq\max\Phi_{\pi}-\alpha$ implies
$\left\lvert\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}})-\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}}^{*})\right\rvert\leq\delta,$ (5.34)
Proof of ˜5.10.
By contradiction, suppose that there exists $\delta>0$ such that for all $\alpha=1/n$ with $n\in\mathbb{N}$, there exists $\widehat{\boldsymbol{y}}_{n}$ with $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{n})\geq\max\Phi_{\pi}-1/n$ and $\left\lvert\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}}_{n})-\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}}^{*})\right\rvert>\delta$. By compactness, take a subsequence and assume that $\widehat{\boldsymbol{y}}_{n}\to\widehat{\boldsymbol{y}}_{\infty}$ as $n\to\infty$. Letting $n\to\infty$, we get that $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{\infty})\geq\max\Phi_{\pi}$ and $\left\lvert\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}}_{\infty})-\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}}^{*})\right\rvert\geq\delta$. This is a contradiction because $\widehat{\boldsymbol{y}}_{\infty}$ would then be a maximizer of the potential, thus a Bayesian Wardrop equilibrium, with a different total payoff. ∎
This concludes the proof of Theorem˜5.9. ∎
6 Optimality of Deterministic BCWE
The following proposition identifies a class of games with binary action sets in which, in order to maximize the designer’s payoff function $\psi$, attention can be restricted to deterministic flows conditional on the state. A BCWE $\mu$ is said to be deterministic if, in each state $\theta\in\Theta$, a single flow profile is implemented with probability one. Hence, a deterministic BCWE can be identified with a mapping $\boldsymbol{y}(\,\cdot\,)\colon\Theta\to\mathcal{Y}$ such that, for all $a,b\in\mathcal{A}$,
$\sum_{\theta\in\Theta}p(\theta)\,y_{a}(\theta)u_{a}(\boldsymbol{y}(\theta),\theta)\geq\sum_{\theta\in\Theta}p(\theta)\,y_{a}(\theta)u_{b}(\boldsymbol{y}(\theta),\theta).$ (6.1)
Under complete information, the notion of deterministic BCWE coincides with that of a Wardrop equilibrium.
Proposition 6.1.
Consider a game where $\mathcal{A}$ contains exactly two actions; for every $a\in\mathcal{A}$ and $\theta\in\Theta$, the mapping $\boldsymbol{y}\mapsto u_{a}(\boldsymbol{y},\theta)$ is convex and the mapping $\boldsymbol{y}\mapsto y_{a}u_{a}(\boldsymbol{y},\theta)$ is concave. Moreover, assume that the designer payoff function $\boldsymbol{y}\mapsto\psi(\boldsymbol{y},\theta)$ is concave for every $\theta\in\Theta$. Then, for every BCWE of $\Gamma$, there exists a deterministic BCWE for which the expected designer payoff is weakly higher.
@@ 6 Optimality of Deterministic BCWE
Consider a BCWE outcome $\mu$. Then
$\begin{split}&\forall a,b\in\mathcal{A},\ \sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\geq\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta).\end{split}$ (6.2)
Since there are two actions, we have $y_{b}=1-y_{a}$ for $b\neq a$. The incentive constraint corresponding to recommendation $a$ and deviation $b\neq a$ can therefore be rewritten as
$\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\geq\sum_{\theta\in\Theta}p(\theta)\int(1-y_{b})u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta),$ (6.3)
or equivalently,
$\begin{split}&\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)+\sum_{\theta\in\Theta}p(\theta)\int y_{b}u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\\ &\quad\geq\sum_{\theta\in\Theta}p(\theta)\int u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta).\end{split}$ (6.4)
For each state $\theta$, define the average flow profile
$\bar{\boldsymbol{y}}(\theta)\coloneqq\int\boldsymbol{y}\ \textup{d}\mu(\boldsymbol{y}\mid\theta),\quad\text{with components }\bar{y}_{a}(\theta)=\int y_{a}\ \textup{d}\mu(\boldsymbol{y}\mid\theta).$ (6.5)
By concavity of $\boldsymbol{y}\mapsto y_{a}u_{a}(\boldsymbol{y},\theta)$ and convexity of $\boldsymbol{y}\mapsto u_{b}(\boldsymbol{y},\theta)$, Jensen’s inequality yields
$\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)$ $\leq\sum_{\theta\in\Theta}p(\theta)\,\bar{y}_{a}(\theta)u_{a}(\bar{\boldsymbol{y}}(\theta),\theta),$ (6.6)
$\sum_{\theta\in\Theta}p(\theta)\int y_{b}u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)$ $\leq\sum_{\theta\in\Theta}p(\theta)\,\bar{y}_{b}(\theta)u_{b}(\bar{\boldsymbol{y}}(\theta),\theta),$ (6.7)
$\sum_{\theta\in\Theta}p(\theta)\int u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)$ $\geq\sum_{\theta\in\Theta}p(\theta)\,u_{b}(\bar{\boldsymbol{y}}(\theta),\theta).$ (6.8)
Thus, Eq.˜6.4 implies that $\theta\mapsto\bar{\boldsymbol{y}}(\theta)$ satisfies the deterministic BCWE incentive constraints. Finally, concavity of $\boldsymbol{y}\mapsto\psi(\boldsymbol{y},\theta)$ implies
$\sum_{\theta\in\Theta}p(\theta)\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\leq\sum_{\theta\in\Theta}p(\theta)\psi(\bar{\boldsymbol{y}}(\theta),\theta),$ (6.9)
which completes the proof. ∎
Appendix A List of Symbols and Acronyms
$\mathcal{A}$ action set, defined in Definition˜2.1